2. What are the units of half life? How many half lives have expired when a radioactive object decays
to 12.5% of its original radioactivity?
Answer #1

Hi, Please find your answer. If you have any query regarding to question feel free to ask.

Answer:

As we know half life is the time required for any quantity to reduce their 50 % of its initial value. The units of half life is Time.

To calculate the half life we use this formula-

N(t) =N(0) (1/2)n

Suppose initial quantity is 100, after a radioactive objective decay the value become 12.5 % of initial

N0 = 100

Nt = 12.5

12.5 = 100 (1/2)n

12.5/100 = (1/2)n

(1/8) = (1/2)n

(1/2)3= (1/2)n

On comparing this n =3

Three half lives have expired to decay 12.5 % from its orignal radioactivity.